PLEASE HELP!
FIND ALL ROOTS OF THE POLYNOMIAL x3-x2+16x-16.

Answer by venugopalramana(3286)

PLEASE HELP!
FIND ALL ROOTS OF THE POLYNOMIAL x3-x2+16x-16.
WE FIND F(1)=0....SO X-1 IS A FACTOR
DIVIDING WITH X-1
1|1-1+16-16
0+1+0 +16
---------------
1+0 +16+0
QUOTIENT = X^2+16 = (X+4i)(X-4i)
HENCE ALL THE ROOTS ARE
X=1,4i,-4i